OpenStudy (anonymous):

Describe the difference between thermal energy and temperature. Use the following words to help structure your answer: particles, kinetic (movement) energy, thermal energy, and measurement

OpenStudy (anonymous):

Hey ;D Welcome to OpenStudy ;)

OpenStudy (anonymous):

Temperature is the average kinetic energy for a sample of something. Kinetic energy of each molecule = (1/2)x m x v^2. So for a sample of water at 300 Kelvin, there molecules that have a higher kinetic energy and there are ones that have a higher energy. Since all water molecules have the same mass, the only difference between one molecule and another is its velocity. Kinetic energy per mole of a substance = (3/2) x universal gas constant x Temperature in Kelvins. So you have: joules/mol = (3/2) x 8.31 joules/(mol*Kelvin) x Kelvins. The units cancel out on the right side to produce the same units on the left side.